### v1

# SmartThinker: Learning to Compress and Preserve Reasoning by Step-Level Length Control

Abstract
Large reasoning models (LRMs) have exhibited remarkable reasoning capabilities through inference-time scaling, but this progress has also introduced considerable redundancy and inefficiency into their reasoning processes, resulting in substantial computational waste. Previous work has attempted to mitigate this issue by penalizing the overall length of generated samples during reinforcement learning (RL), with the goal of encouraging a more concise chains of thought. However, we observe that such global length penalty often lead to excessive compression of critical reasoning steps while preserving unnecessary details in simpler ones, yielding a suboptimal trade-off between accuracy and efficiency. To address this issue, we propose SmartThinker, a two-stage learnable framework designed to enable fine-grained control over the length of reasoning chains based on the importance of each individual step. In the first stage, SmartThinker adapts a reasoning model to a short-form reasoning mode through rejection sampling combined with supervised fine-tuning (SFT). In the second stage, SmartThinker applies Step-Level Length Control Policy Optimization (SCPO) to refine the model output distribution, which increases the proportion of length allocated to critical steps while reducing redundancy in less important ones. SCPO consists of four core components: an online importance estimator, a step-level length control reward function, a step-level generalized advantage estimation (S-GAE) and a difficulty-adaptive clipping strategy. Working in concert, these components enable SCPO to implement differentiated length control across reasoning steps. Empirical results across multiple reasoning benchmarks and various backbone models demonstrate that SmartThinker significantly reduces redundant reasoning while achieving comparable or even superior performance to existing methods.
(a) Output distribution of the model trained using overall length penalty, and the ratio of effective step length to total length.
(b) Output distribution of models trained using SmartThinker, and the ratio of effective step length to total length.
Figure 1: Comparison of the output distributions between the model trained with global length penalty and that trained with SmartThinker on problems of varying difficulty. Please see Appendix A.3 for experimental details.
Introduction
LRMs have demonstrated the effectiveness of inference-time scaling, during which they often generate extensive chains of thought involving self-reflection and verification (Guo et al. 2025; Jaech et al. 2024). Although this approach substantially improves performance across many domains, such as mathematical problem solving (Yang et al. 2024; Luo et al. 2024) and agentic assistants (Chen et al. 2025a; Wu, Zhu, and Liu 2025), it also faces the significant challenge of overthinking (Chen et al. 2024; Cuadron et al. 2025), in which the model generates excessive and inefficient reasoning even for questions with straightforward solutions. (Feng et al. 2025; Kumar et al. 2025). This results in redundant token generation and considerable computational waste (Sui et al. 2025; Qu et al. 2025).
To address this issue, previous studies have attempted to penalize the overall length of reasoning chains within a RL framework to encourage the generation of more concise reasoning chains (Aggarwal and Welleck 2025; Arora and Zanette 2025; Hou et al. 2025; Zhang and Zuo 2025). However, these approaches typically focus only on the overall length of the final output and lack fine-grained feedback on intermediate reasoning steps. Under such a sparse reward signals, models are prone to reward hacking, where the model compresses critical reasoning steps too aggressively while retaining unnecessary details in simpler steps, ultimately compromising both reasoning efficiency and quality. As shown in Fig. 1(a), we visualize the relationship between the importance and length of each reasoning step to better understand the impact of this overall length penalty mechanism. The results reveal that regardless of whether the problem is simple or complex, both the number and total length of high-importance steps remain quite limited, displaying a sharply skewed distribution. These findings indicate that the global length penalty mechanism compresses step lengths indiscriminately and fails to improve the quality of model reasoning effectively. This observation raises a key question: Can we apply differentiated length control at the step level to achieve a better balance between reasoning performance and computational efficiency?
In this paper, we propose SmartThinker, a two-stage learnable framework designed to enable fine-grained control over the length of reasoning chains based on step-level importance. In the first stage, SmartThinker leverages rejection sampling and SFT to quickly adapt a reasoning model to a short-reasoning mode. Although this process may lead to a slight performance drop, it significantly accelerates convergence in the subsequent RL phase. In the second stage, SmartThinker introduces Step-Level Length Control Policy Optimization (SCPO) to perform fine-grained length control across individual reasoning steps. The core idea of SCPO is to leverage differences in importance across reasoning steps to guide the model to impose stronger penalty on redundant steps while relaxing length constraints on critical ones, thereby achieving a more effective balance between reasoning quality and computational efficiency. SCPO comprises four key components: (i) Online importance estimator: This module evaluates the importance score of each reasoning step in real time by measuring the change in the probability of generating the final answer when the step is removed. This approach eliminates the need for training complex reward models or relying on additional human annotations. Moreover, if a reasoning step contains keywords associated with reasoning transitions (e.g., ”but,” ”however”), an additional scoring mechanism is triggered. This supplementary score encourages the model to engage in self-reflection and verification when handling challenging problems, improving its ability to explore correct answers effectively. (ii) Step-level length control reward function: This component dynamically adjusts the length penalty coefficient based on both the step importance and the problem difficulty, providing differentiated reward signals for each step. (iii) Step-level generalized advantage estimation (S-GAE): S-GAE computes the cumulative discounted advantage return for each step, quantifying its long-term contribution to subsequent reasoning, which effectively mitigates the length bias introduced by direct accumulation in the original Group Relative Policy Optimization (GRPO) (Shao et al. 2024). (iv) Difficulty-adaptive clipping strategy: This strategy dynamically adjusts the upper and lower bounds of the clipping based on the problem difficulty, assigning wider exploration ranges to more challenging problems while maintaining stability on easier ones. The coordinated interaction of these four components enables SCPO to achieve differentiated length control across individual reasoning steps.
Extensive empirical evaluations show that SmartThinker significantly reduces redundant reasoning steps while achieving performance comparable or even superior to existing methods. In summary, our contributions are as follows:
  • We propose SmartThinker, a two-stage learnable framework that enables fine-grained control over the length of reasoning chains based on step-level importance, which effectively mitigates reward hacking caused by global length penalty and achieves a better trade-off between reasoning accuracy and computational efficiency.
  • We introduce SCPO, which leverages step-level importance to generate dense process reward signals, guiding the model to penalize the length of redundant steps more strongly while relaxing constraints on critical ones.
  • We conduct extensive experiments to evaluate the effectiveness of SmartThinker in reducing redundant reasoning, and detailed ablation studies further confirm the contribution of each component within the framework.
Related Work
Efficient Reasoning Models
Reasoning models improve their ability to solve complex problems by generating extended chains of thought (Ye et al. 2025; Chen et al. 2025b). However, excessively long reasoning chains often introduce redundancy and waste computational resources (Kumar et al. 2025; Fu et al. 2024). To address this, prior work has focused on enhancing reasoning efficiency without sacrificing accuracy (Sui et al. 2025; Qu et al. 2025). Existing methods mainly fall into two categories: SFT-based approaches and RL. SFT-based methods either create synthetic datasets with shorter reasoning chains to train models for concise reasoning (Xia et al. 2025; Han et al. 2024; Xu et al. 2025) or teach models to dynamically compress reasoning during inference (Zhang et al. 2025b; Yan et al. 2025). Meanwhile, RL has emerged as a popular strategy, designing reward functions that penalize longer chains to encourage brevity (Aggarwal and Welleck 2025; Arora and Zanette 2025; Hou et al. 2025). However, simply penalizing overall chain length overlooks the varying importance of individual reasoning steps, often causing models to eliminate crucial steps and compromise reasoning quality. In contrast, SmartThinker explicitly models the importance of different steps, imposing stronger penalties on redundant ones while relaxing constraints on critical steps, thus achieving a better balance between efficiency and accuracy.
Hybrid Reasoning Models
[…]
Figure 2: Overview of the proposed SmartThinker.
Method
In this section, we present an overview of the proposed SmartThinker. As shown in Fig. 2, SmartThinker consists of two stages: (i) Short-Reasoning Mode Warm-up, where we adapt a reasoning model to a short-reasoning mode by rejection sampling and SFT, and (ii) RL with SCPO, where the model is trained to impose stronger penalty on redundant steps while relaxing length constraints on critical ones.
Preliminaries
GRPO (Shao et al. 2024) eliminates the need for a value function by estimating advantages in a group-relative manner, thereby significantly reducing memory consumption. Given a specific question-answer (QA) pair $(q,a)$, the behavior policy $\pi_{\theta_{\text{old}}}$ samples a group of $G$ individual responses $\{o_{i}\}_{i=1}^{G}$. The advantage of the $i$-th response is then computed by normalizing the group-level rewards $\{R_{i}\}_{i=1}^{G}$:
${A}_{i,t}=\frac{r_{i}-\text{mean}(\{R_{i}\}_{i=1}^{G})}{\text{std}(\{R_{i}\}_{i=1}^{G})}$ (1)
GRPO subsequently optimizes a clipped objective augmented with a directly imposed KL penalty term:
$\mathcal{J}(\theta)=\mathbb{E}[($ $\min(r_{i,t}(\theta){A}_{i,t},\,\text{clip}(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon){A}_{i,t})$
$-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}}))]$ (2)
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ (3)
Short-Reasoning Mode Warm-up
Directly fine-tuning a long-reasoning model into a short-reasoning model via RL often leads to unstable training and slow convergence, as the model initially spends a significant amount of time generating overly long responses and needs multiple rounds of feedback to discover suitable reasoning paths. To mitigate this, we combine rejection sampling with SFT to help the model quickly adapt to a short-reasoning mode before starting RL training, providing a better initialization. Specifically, for each question in a given corpus, we use the current reasoning model to generate multiple candidate responses and select the shortest correct one to build a synthetic dataset. We then perform SFT on this dataset to align the model more closely with the desired short-reasoning behavior. Because the training data are generated by the model itself, their distribution closely matches the model’s own output, helping it adapt to short reasoning without significant performance degradation.
SCPO
Online Importance Evaluator
The importance of a reasoning step depends on whether it increases the probability of generating the correct answer. Based on this insight, we quantify the importance score of a step as the change in the probability of producing the correct answer when the step is removed. When given a QA pair $(q,a)$, SCPO first samples G responses $\{o_{i}\}_{i=1}^{G}$. The $o_{i}$ consists of a reasoning trajectory $S_{i}=\left\{s_{i1},s_{i2},...,s_{ik_{i}}\right\}$ and a predicted answer $a_{i}$, where $s_{i,j}$ denotes the $j$-th reasoning step of $o_{i}$ and $k_{i}$ is the total number of steps in $o_{i}$. The importance score of the $j$-th reasoning step of $o_{i}$ is expressed as:
$d_{i,j}=\left\{\begin{matrix}0,&\text{if}\ p_{1}\leq p_{2},\\ \frac{p_{1}^{2}-p_{2}^{2}}{p_{1}^{2}\cdot l_{i,j}},&\text{if}\ p_{1}>p_{2}.\end{matrix}\right.$ (4)
where $p_{1}=P(a\mid q,S_{i})$ is the probability of generating the correct answer using the complete reasoning trajectory, $p_{2}=P(a\mid q,S_{i}\setminus\left\{s_{i,j}\right\})$ is the probability after removing the step $s_{i,j}$ and $l_{i,j}$ is the length of the $j$-th reasoning step.
As shown in Eq.4, if removing a reasoning step does not reduce the probability of generating the correct answer, the step is deemed unimportant and assigned an importance score of zero. Otherwise, the score is computed using the relative squared difference to emphasize the variation in importance across steps. The score is then normalized by step length to reduce bias from varying step lengths.
However, we find that relying solely on the above importance scores to guide model updates may suppress complex reasoning behaviors such as self-reflection and verification, leading to suboptimal performance on more challenging problems. To address this issue, we assign additional importance scores to steps containing keywords associated with reasoning transitions (e.g., ”but,” ”however”). These supplementary scores increase with problem difficulty, encouraging the model to engage in deeper reflection and verification when handling complex problems, while maintaining concise reasoning for simpler ones. The final importance score of the $j$-th reasoning step of $o_{i}$ is expressed as:
[…]
where
[…]
$\rho=1-\frac{N_{c}}{N}$ (7)
denotes the indicator function and problem difficulty, respectively. Here, $N_{c}$ is the number of correct responses and $N$ is the total number of sampled responses.
Step-Level Length Control Reward Function
An efficient and intelligent reasoning model should follow two core principles when generating reasoning processes. First, critical reasoning steps should be given sufficient length to support inference-time scaling, while auxiliary steps should remain concise to avoid redundancy. Second, the number of steps should adapt dynamically to problem difficulty: complex problems demand more steps for solution quality, whereas simpler ones benefit from brevity. Thus, the reward function must capture two key aspects: aligning step length with each step’s importance and adjusting the overall number of steps based on problem difficulty.
To achieve the above objectives, we design a step-level length control reward function. The reward of the $j$-th reasoning step of $o_{i}$ is expressed as:
$r_{i,j}=\left\{\begin{matrix}(1-k_{1}\sigma(\widetilde{l}_{i,j}))(1-k_{2}\sigma(\widetilde{n}_{i})),&\text{if }a_{i}=a,\\ -e^{-\frac{\rho\cdot\widetilde{d}_{i,j}^{\prime}}{k_{0}}},&\text{if }a_{i}\neq a.\end{matrix}\right.$ (8)
where
$k_{1}$ $=k_{0}\cdot(1-\widetilde{d}_{i,j}^{\prime})$ (9)
$k_{2}$ $=k_{0}\cdot(1-\rho)$ (10)
represent the step length penalty coefficient and step number penalty coefficient, respectively, and $k_{0}$ is the base penalty coefficient. Here,
$\widetilde{l}_{i,j}=\frac{l_{i,j}-\underset{\left\{i,j|a_{i}=a\right\}}{\text{mean}}(l_{i,j})}{\underset{\left\{i,j|a_{i}=a\right\}}{\text{std}}(l_{i,j})}$ (11)
$\widetilde{n}_{i}=\frac{n_{i}-\underset{\left\{i,j|a_{i}=a\right\}}{\text{mean}}(n_{i})}{\underset{\left\{i,j|a_{i}=a\right\}}{\text{std}}(n_{i})}$ (12)
denotes the standardized step length and step number for the correct responses, respectively. To ensure that the importance scores fall within the range $\left[0,1\right]$ and to prevent mutual interference between positive and negative samples, we apply min-max normalization separately to each group:
$\widetilde{d}_{i,j}^{\prime}=\left\{\begin{matrix}\displaystyle\frac{\widetilde{d}_{i,j}-\min(\widetilde{d}_{i,j}^{+})}{\max(\widetilde{d}_{i,j}^{+})-\min(\widetilde{d}_{i,j}^{+})},&\text{if }a_{i}=a\\ \displaystyle\frac{\widetilde{d}_{i,j}-\min(\widetilde{d}_{i,j}^{-})}{\max(\widetilde{d}_{i,j}^{-})-\min(\widetilde{d}_{i,j}^{-})},&\text{if }a_{i}\neq a\end{matrix}\right.$ (13)
where $\widetilde{d}_{i,j}^{+}=\left\{\widetilde{d}_{i,j}\mid a_{i}=a\right\}$ and $\widetilde{d}_{i,j}^{-}=\left\{\widetilde{d}_{i,j}\mid a_{i}\neq a\right\}$ represent the set of importance scores corresponding to correct and incorrect answers, respectively.
In formulation design, we do not impose length constraints on incorrect answers, allowing them to expand freely to explore potential correct solutions. Meanwhile, the reward for incorrect responses diminishes as problem difficulty decreases, meaning that incorrect responses to easier questions are penalized more heavily, thereby reinforcing the model’s decision boundaries in scenarios where high accuracy is expected.
S-GAE
In the original GRPO (Shao et al. 2024) with process supervision, the advantage function is estimated by summing the normalized future rewards of subsequent steps:
$A_{i,j}=\sum_{n=0}^{k_{i}-j}\widetilde{r}_{i,j+n}$ (14)
where
$\widetilde{r}_{i,j}=\frac{r_{i,j}-\underset{i,j}{\mathrm{mean}}(r_{i,j})}{\underset{i,j}{\mathrm{std}}(r_{i,j})}$ (15)
denotes the normalized rewards.
However, this approach introduces a length bias, as correct responses with more reasoning steps accumulate larger advantage values, prompting the model to generate unnecessarily long reasoning sequences. To address this, S-GAE applies a discount factor $\gamma$ to downweight the influence of distant steps in the advantage calculation, thereby reducing the bias between long and short sequences. The S-GAE formula is given below:
$A_{i,j}=\sum_{n=0}^{k_{i}-j}\gamma^{n}\cdot\widetilde{r}_{i,j+n}$ (16)
Finally, S-GAE distributes the step-level advantage uniformly across all tokens within the step, which fundamentally distinguishes it from outcome-reward-based advantage estimation. In outcome-reward-based advantage estimation, each token is uniformly assigned a normalized final return as its advantage, without differentiating between the varying contributions of critical and redundant steps during policy optimization. This coarse-grained advantage estimation results in suboptimal and inefficient policy updates, especially for complex reasoning tasks. In contrast, S-GAE provides dense process reward signals and computes differentiated advantage estimates for each step, significantly improving the stability and accuracy of the advantage estimation.
Difficulty-Adaptive Clipping Strategy
Clip-Higher, first proposed in DAPO (Yu et al. 2025), mitigates entropy collapse by expanding the clipping range, allowing low-probability tokens more room for exploration. However, in simpler problems, the model often generates correct answers with high confidence, and low-probability tokens frequently lead to incorrect reasoning. In such cases, a large $\varepsilon_{\text{high}}$ can steer the model away from the correct path. Conversely, a smaller $\varepsilon_{\text{high}}$ improves stability on simpler tasks but may limit exploration in more complex ones. To balance exploration and exploitation across varying problem difficulties, we propose a difficulty-adaptive clipping strategy:
$\varepsilon_{\text{low}}$ $=\varepsilon-\delta_{1}\cdot(1-\rho)$ (17)
$\varepsilon_{\text{high}}$ $=\varepsilon+\delta_{2}\cdot\rho$ (18)
where $\delta_{1}$ and $\delta_{2}$ represent the magnitudes of decrease and increase, respectively.
As shown in Eq.17 and Eq.18, for low-difficulty problems, smaller values of $\varepsilon_{\text{high}}$ and $\varepsilon_{\text{low}}$ limit the upward adjustment of low-probability tokens and the downward adjustment of high-probability tokens. This curbs excessive exploration and encourages the model to focus on known correct reasoning paths. In contrast, for high-difficulty problems, moderately increasing $\varepsilon_{\text{high}}$ and $\varepsilon_{\text{low}}$ loosens these bounds, fostering exploration of potentially correct reasoning paths and improving performance on more complex tasks.
Table 1: Pass@k, Maj@k and AvgLen performance on various math reasoning benchmarks
Models AIME25 AIME24 AMC23 MinervaMATH MATH Olympiad-Bench Avg.
Pass@k (%)
DeepSeek-R1-1.5B 36.7 50.0 90.0 40.0 91.8 55.2 60.6
Alpha-0.1 40.0(+3.3) 60.0(+10.0) 95.0(+5.0) 43.3(+3.3) 92.2(+0.4) 60.0(+4.8) 65.1(+4.5)
L1-Max 40.0(+3.3) 50.0(+0.0) 95.0(+5.0) 43.3(+3.3) 92.4(+0.6) 58.6(+3.4) 63.2(+2.6)
Thinkless 43.3(+6.6) 56.7(+6.7) 90.0(+0.0) 42.2(+2.2) 91.2(-0.6) 61.6(+6.4) 64.1(+3.5)
[…]
Maj@k (%)
DeepSeek-R1-1.5B 26.7 43.3 77.5 31.3 87.2 45.2 51.8
Alpha-0.1 30.0(+3.3) 33.3(-10.0) 85.0(+7.5) 31.6(+0.3) 86.6(-0.6) 49.8(+4.6) 52.7(+0.9)
L1-Max 33.3(+6.6) 30.0(-13.3) 87.5(+10.0) 34.6(+3.3) 88.8(+1.6) 52.0(+6.8) 54.3(+2.5)
Thinkless 26.7(+0.0) 40.0(-3.3) 85.0(+7.5) 29.7(-1.6) 83.2(-4.0) 50.6(+5.4) 52.5(+0.7)
[…]
AvgLen
DeepSeek-R1-1.5B 7307 7380 5331 4798 3809 5953 5763
Alpha-0.1 6345(86%) 6696(90%) 3802(71%) 2511(52%) 2135(56%) 4420(74%) 4318(75%)
L1-Max 3526(48%) 3557(48%) 3255(61%) 3325(69%) 3023(79%) 3409(57%) 3349(58%)
Thinkless 6303(86%) 6287(85%) 3637(68%) 2804(58%) 1859(48%) 3997(67%) 4147(72%)
[…]
Experiments
Models and Datasets
We use DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B as base models and train these models on a randomly sampled subset of 3000 QA pairs from the DeepScaleR-Preview dataset (Luo et al. 2025b). For evaluation, we use test sets from six diverse reasoning benchmarks: AIME24, AIME25 (Veeraboina 2023), AMC23, MATH (Lewkowycz et al. 2022), MinervaMATH (Hendrycks et al. 2021), and Olympiad-Bench (Gao et al. 2024). Please see Appendix A.1 for detailed descriptions of these benchmarks.
We evaluate performance using three main metrics: (i) Pass@k, the probability that at least one of the top $k$ samples is correct under a given decoding strategy, reflecting the upper bound of the model’s reasoning ability. (ii) Maj@k, the accuracy obtained through majority voting across $k$ samples, which measures the stability of the model’s output. (iii) AvgLen, the average length of all sampled responses, indicating the model’s reasoning efficiency. For AIME24, AIME25, and AMC23, we set $k=10$ due to the small size of their test sets. For all other datasets, we use $k=5$.
Note that we do not use Pass@1 as our primary evaluation metric, since improvements in Pass@1 mainly reflect better output ranking (Zhu et al. 2025). In other words, the model becomes more skilled at placing the correct answer first rather than solving previously unsolvable problems. In contrast, Pass@k (for $k>1$) allows multiple attempts, allowing the model to explore potentially correct answers even if the first attempt fails. This setting better reflects the model’s true reasoning ability when sufficient exploration is allowed.
Baselines
The baselines we select can be grouped into two categories: (i) models based on global length penalty, including Alpha-0.1 (Arora and Zanette 2025) and L1-Max (Aggarwal and Welleck 2025), and (ii) models that incorporate both long- and short-reasoning modes, such as Thinkless (Fang, Ma, and Wang 2025) and AdaptThink (Zhang et al. 2025a).
Experiment Settings
The hyperparameters are configured as $k_{0}=0.6$, $\gamma=0.95$, $\delta_{1}=0.03$, and $\delta_{2}=0.08$. Due to space limitation, we comprehensively introduce our experimental setting in Appendix A.4.
Main Results
Output Distribution Analysis
As shown in Fig. 1(b), SmartThinker yields a flatter output distribution than models using overall length penalty. The proportion of effective steps increases by 15.5% for low-difficulty questions and by 9.3% for high-difficulty questions, indicating a reduction in redundant steps and a greater focus on critical reasoning. Moreover, the proportion of the total length attributed to effective steps has increased by approximately 10%. This improvement indicates that SmartThinker successfully relaxes length constraints for critical steps while imposing stricter length penalty on redundant ones, refining the model’s output distribution and enabling higher-quality reasoning. We present case studies of different baselines in Appendix A.6 to illustrate SmartThinker’s effectiveness.
Figure 3: Pass@5 accuracy and AvgLen at different training steps on the MATH dataset.
Figure 4: Effect of additional scores on reasoning behavior and Pass@k performance across different difficulty levels.
Figure 5: Entropy trends driven by problem difficulty during training. Since each batch of samples is used to update the policy model over four iterations, both the entropy and problem difficulty are averaged every four steps.
Effectiveness and Efficiency Analysis
As shown in Table 1, SmartThinker achieves the highest reasoning accuracy while using the fewest tokens. For example, compared to Alpha-0.1, SmartThinker improves the upper bound of reasoning performance on the challenging AIME24 dataset, increasing Pass@10 accuracy by 3.3% while reducing average token usage by about 1.5k. On AIME25, SmartThinker also increases Maj@10 accuracy by 3.3%, reflecting a more consistent ability to produce correct answers. These improvements stem from SmartThinker’s fine-grained control over the reasoning process, enabling a better balance between performance and computational efficiency. Notably, L1-Max, despite being trained on the more powerful DeepScaleR (Luo et al. 2025b) model, achieves only 50% Pass@10 and 30% Maj@10 on AIME2024, suggesting that its global length penalty may excessively compress critical reasoning steps, sacrificing accuracy for token savings. Moreover, Thinkless has limited improvements in reasoning accuracy on simpler datasets like AMC23 and MATH, possibly because switching to shorter reasoning hampers sufficient exploration of correct answers. In summary, these results confirm the superiority of SmartThinker, which consistently delivers more accurate reasoning with greater computational efficiency. Please refer to Appendix A.5 for the experimental results of the 7B model.
Figure 6: Hyper Parameter Selection
Ablation Study
Impact of SFT on RL
As shown in Fig. 3, our method consistently boosts reasoning capability while reducing token usage, regardless of whether SFT is applied. Although the introduction of SFT leads to a slight drop in accuracy, it effectively guides the model to adapt to a shorter reasoning pattern, significantly accelerating convergence during the subsequent RL stage. For example, without SFT, the model tends to generate lengthy responses, requiring 122 hours to complete 1,000 training steps. In contrast, with SFT, the model completes 1,300 steps in just 68 hours, achieving comparable accuracy while substantially improving training efficiency.
Effectiveness of Additional Scores
Fig. 4 illustrates the impact of incorporating additional scores on both the model’s reasoning behavior and performance. As shown in Fig. 4a, without the use of additional scores, the model significantly reduces the use of reasoning transition words, which impairs its ability to explore alternative solutions and leads to a drop in accuracy, as reflected in Fig. 4b. This effect is particularly pronounced on easier problems. The underlying reason is that the model often generates multiple correct solutions for simple problems and reaches the answer early in the reasoning process. As a result, subsequent reflection and verification steps are considered low-importance and become weakened or even discarded during training, pushing the model toward shallow, non-reflective reasoning strategies. In contrast, when additional scores are used to selectively compensate for the importance of specific steps, the model achieves a better balance between concise reasoning and sufficient exploration, which leads to an overall improvement in the upper bound of reasoning performance.
Entropy Curve Analysis
As shown in Fig. 5, higher problem difficulty correlates with increased output entropy, indicating greater exploration, while lower difficulty leads to reduced entropy and more deterministic reasoning. This balance between exploration and exploitation is achieved through our difficulty-adaptive clipping strategy, which adjusts the clipping range based on problem difficulty to stabilize model updates and prevent entropy collapse or explosion.
Hyper Parameter Selection
Effect of $\delta_{1}$ on Entropy
Following DAPO (Yu et al. 2025), we set $\delta_{2}$ to 0.08 and compare the effects of a larger $\delta_{1}$ with the smaller value used in this work. As shown in Fig. 6(a), a larger $\delta_{1}$ raises entropy but eventually causes entropy collapse, as excessive exploration during moderately complex tasks drives the model off the correct reasoning path. As training continues, this drift escalates, destabilizing training. In contrast, a smaller $\delta_{1}$ curbs entropy growth, maintaining a healthier balance between exploration and exploitation across problem difficulties.
Impact of $\gamma$ on Reasoning Performance
As shown in Fig. 6(b), when $\gamma=1$, the model tends to generate excessively long sequences. As $\gamma$ decreases, the average sequence length also shortens. However, setting $\gamma$ too low greatly reduces the contribution of later steps to the overall advantage, leading to less accurate advantage estimates and potentially harming model performance. Balancing these trade-offs, we set $\gamma=0.95$, which effectively mitigates length bias while maintaining stable performance.
Stability Analysis of $k_{0}$
As shown in Fig. 6(c), we visualize how different values of $k_{0}$ affect the model’s reasoning accuracy and efficiency. When $k_{0}=0.8$, the model uses the fewest tokens but suffers a notable drop in accuracy. This happens because a higher $k_{0}$ pushes the model to reduce token usage more aggressively for higher rewards, leaving too few tokens to adequately explore the reasoning space, thus lowering accuracy. As $k_{0}$ decreases, token usage rises, and accuracy first improves but then declines. This occurs because smaller $k_{0}$ values narrow the feasible ranges for the step length penalty coefficient $k_{1}$ and the step number penalty coefficient $k_{2}$, making it harder to capture differences in the importance of individual reasoning steps when calculating the reward for correct answers. Consequently, the model focuses on compressing the reasoning length at the expense of reasoning quality, ultimately reducing accuracy. Considering both reasoning accuracy and efficiency, we choose $k_{0}=0.6$ as it strikes a better balance between performance and computational cost.
Conclusion
In this paper, we propose SmartThinker, a two-stage learnable framework that enables fine-grained control over the length of reasoning chains based on step-level importance. It begins with a warm-up phase for short-reasoning modes, followed by Step-Level Length Control Policy Optimization (SCPO). The warm-up phase provides a better initialization for subsequent RL, which speeds up convergence. SCPO exploits differences in step importance to penalize redundant steps while relaxing length constraints for critical ones. Extensive experiments on various math reasoning benchmarks demonstrate that SmartThinker improves both reasoning effectiveness and efficiency.
Appendix A A Appendix
A.1 Benchmark Datasets
In this section, we will describe in detail all the datasets used in our experiment.
  • AIME24–25 comprises 30 questions from the 2024 and 2025 American Invitational Mathematics Examination, with 15 fill-in-the-blank questions per exam. These questions are more difficult than AMC and span number theory, combinatorics, geometry, and algebra.
  • AMC10/12 consists of 25 multiple-choice questions each for the AMC10 (up to 10th grade) and AMC12 (up to 12th grade). Each competition consists of 25 multiple-choice questions, totaling 975 questions across 39 tests. Questions progress from basic algebra and geometry to introductory probability and counting, covering various tasks for LLM reasoning evaluation.
  • MinervaMATH comprises 12,500 high-school–level contest questions. Each includes detailed solution steps and spans prealgebra through precalculus.
  • MATH is a challenging dataset used to evaluate models’ complex multi-step reasoning abilities in mathematics, featuring problems that require precise logical reasoning where a single error can invalidate an entire solution. It has automatically checkable answers, allowing for outcome supervision without human input, and is employed in comparing outcome and process supervision methods.
  • OlympiadBench is an Olympiad-level bilingual multimodal scientific benchmark, featuring 8,476 problems from Olympiad-level mathematics and physics competitions, including the Chinese college entrance exam. Each problem is detailed with expert-level annotations for step-by-step reasoning. Notably, the best-performing model, GPT-4V, attains an average score of 17.97% on OlympiadBench, with a mere 10.74% in physics, highlighting the benchmark rigor and the intricacy of physical reasoning.
Table 2: Pass@k, Maj@k and AvgLen performance on various math reasoning benchmarks
Models AIME25 AIME24 AMC23 MinervaMATH MATH Olympiad-Bench Avg.
Pass@k (%)
DeepSeek-R1-7B 46.7 73.3 95.0 53.3 94.6 62.8 70.9
Alpha-0.1 53.3(+6.6) 70.0(-3.3) 92.5(-2.5) 52.5(-0.8) 94.4(-0.2) 64.0(+1.2) 71.1(+0.2)
AdaptThink 53.3(+6.6) 73.3(+0.0) 92.5(-2.5) 51.8(-1.5) 95.6(+1.0) 66.2(+3.4) 72.0(+1.1)
[…]
Maj@k (%)
DeepSeek-R1-7B 36.7 50.0 90.0 41.5 92.0 56.2 61.1
Alpha-0.1 40.0(+3.3) 56.7(+6.7) 90.0(+0.0) 45.2(+3.7) 92.2(+0.2) 57.4(+1.2) 63.5(+2.4)
AdaptThink 40.0(+3.3) 63.3(+13.3) 92.5(+2.5) 39.3(-2.2) 93.6(+1.6) 59.6(+3.4) 64.7(+3.6)
[…]
AvgLen
DeepSeek-R1-1.5B 6896 6919 4580 4035 3177 5348 5159
Alpha-0.1 6372(92%) 6279(91%) 3812(83%) 2617(65%) 2219(70%) 4537(85%) 4306(83%)
AdaptThink 6258(91%) 5935(86%) 3377(74%) 2177(54%) 1535(48%) 4017(75%) 3883(75%)
[…]
A.2 Baselines
In this section, we will describe in detail all the baselines used in our experiment.
  • Alpha-0.1 introduces a global length penalty mechanism into the PPO algorithm to encourage the model to produce more concise reasoning processes. This approach is fully fine-tuned on both the DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B models, achieving a substantial reduction in redundant reasoning while maintaining high accuracy.
  • L1-Max employs a Length Controlled Policy Optimization (LCPO) mechanism to precisely regulate the length of chain-of-thought reasoning. The method first trains the DeepScaleR-1.5B-Preview model with LCPO, enabling it to follow user-specified token budgets and yielding the base model L1-Exact. Building on this foundation, L1-Max is further trained to generate responses that do not exceed the target length.
[…]
  • AdaptThink leverages RL to encourage the model to skip the reasoning process whenever possible, while incorporating importance sampling techniques to balance the proportion of reasoning and non-reasoning samples during training. The final AdaptThink model is trained on DeepSeek-R1-Distill-Qwen-7B.
A.3 Experimental Details in Fig. 1
We randomly sample 20 questions from the DeepScaleR-Preview-Dataset, generating 8 responses for each question. Problem difficulty is computed using Eq. 7, where questions with a difficulty score less than or equal to 0.5 are categorized as low-difficulty, and those with scores greater than 0.5 are considered high-difficulty. Step importance is calculated using Eq. 4 and subsequently normalized via min-max scaling. Steps with normalized importance scores greater than 0.01 are defined as effective steps. We then compute both the proportion of effective steps and the proportion of total length attributed to these effective steps.
A.4 Experimental Setting Details
In the first stage, we sample 5 responses per training question and retain only the correct and shortest ones, discarding samples exceeding 4096 tokens. This yields around 2000 samples per model for constructing the synthetic dataset. Sampling uses a temperature of 1 and top-p of 0.95. During SFT, we set the batch size to 1 with gradient accumulation of 8. The learning rate is $1\times 10^{-5}$, following a cosine schedule with a warm-up ratio of 0.1, and training runs for three epochs. In the second stage, we train using the TRL framework with a batch size of 8, generating 8 rollouts per question for SCPO reward calculation. We omit the KL penalty, as prior studies show it suppresses exploration (Yu et al. 2025; Liu et al. 2025). The maximum context length is set to 4K tokens during training and extended to 8K for evaluation. Training uses ZeRO Stage 2 (Rajbhandari et al. 2020) on two A6000 GPUs. We use a temperature of 1 and top-p of 0.95 during training, and temperature 0.6 with top-p of 1 during evaluation. To maximize data usage, each batch updates the policy model four times. Hyperparameters are configured as $k_{0}=0.6$, $\gamma=0.95$, $\delta_{1}=0.03$, and $\delta_{2}=0.08$. Due to hardware limits, For the 1.5B model, we perform full fine-tuning, while for the 7B model, we adopt LoRA (Hu et al. 2022) for lightweight tuning. Optimization is performed with the AdamW (Loshchilov and Hutter 2017) optimizer, using $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a learning rate of $1\times 10^{-6}$, following a cosine schedule with a 60-step warm-up.
A.5 Experimental results of the 7B model
Table 2 shows that SmartThinker achieves the best performance in terms of both Pass@k and Maj@k accuracy. On the challenging AIME24-25 dataset, SmartThinker improves accuracy by at least 10% compared to the baseline model. This result demonstrates that our proposed method effectively expands the upper bound of the model in reasoning capabilities and improves the stability of generating correct answers. In contrast, Alpha-0.1 suffers a decline in Pass@k performance on the AIME24 dataset due to its global length penalty mechanism, which indiscriminately compresses the reasoning process. Although AdaptThink minimizes token usage by dynamically switching between short- and long-reasoning modes, it does not fundamentally optimize the model’s output distribution, thus contributing only marginally to performance improvements. Notably, SmartThinker relies solely on LoRA fine-tuning, whereas the other baselines are fine-tuned using full-parameter updates. Overall, SmartThinker achieves a superior trade-off between reasoning efficiency and quality through fine-grained control over the length of reasoning chains.
A.6 Case Studies
We conduct case studies on different baseline models to provide an intuitive demonstration of SmartThinker’s effectiveness. As shown in Fig. 7, for a question that allows a direct answer, the original model generates a lengthy textual reasoning and continues to elaborate even after arriving at the correct answer. This unnecessarily complicates simple problems and results in substantial token redundancy. Although Alpha-0.1 mitigates this redundancy to some extent, it still repeats portions of the reasoning process when producing the final answer. Although Thinkless is capable of switching to short-reasoning modes, its reasoning remains relatively complex. In contrast, SmartThinker provides the final answer directly with a single concise line of reasoning, entirely avoiding such redundancy.
As shown in Fig. 8, when faced with more challenging questions, the original model often engages in extensive purely textual reasoning, followed by multiple rounds of reflection and verification. This significantly increases computational resources required for reasoning without contributing meaningfully to producing the final answer. Although Alpha-0.1 and Thinkless are able to arrive at the correct answers, they still perform redundant self-reflection and verification afterward, leading to unnecessary token consumption. In contrast, SmartThinker relies more on concise and effective statements interwoven with mathematical formulas, outputting only the mathematical derivations and essential textual explanations that are genuinely helpful for producing the final answer. This enables a more efficient reasoning process overall.
Figure 7: Examples of model responses to questions with straightforward solutions across different baselines.
Figure 8: Examples of model responses to challenging questions across different baselines.